OpenStudy (anonymous):

help me triple plz ill give medal check my answeres 3 @sleepyjess

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

cheack if answer is correct plz

OpenStudy (sleepyjess):

this is going to be our equation s+l=11 l=4s-4

OpenStudy (anonymous):

so then wat

OpenStudy (anonymous):

so the answer is 11 4

OpenStudy (sleepyjess):

no

OpenStudy (sleepyjess):

both pieces have to add up to 11

OpenStudy (anonymous):

o ok so then ill get the answer

OpenStudy (anonymous):

so i was right 22 tho right

OpenStudy (sleepyjess):

you have the 9 correct

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

so wats the shorter piace answer

OpenStudy (sleepyjess):

wait no it should be 3 and 8 3+8=11 8=4*3-4 8=12-4 8=8

OpenStudy (anonymous):

so the shorter piece is 3 and the longer piece is 8 right

OpenStudy (sleepyjess):

correct

OpenStudy (anonymous):

k i have 1 more like this quesion to see if its correct may i ask